OpenStudy (anonymous):

which is greater 17/25 or 68/100

OpenStudy (anonymous):

we have 2 values.... 1st one = 17/25 2nd one = 68/25 = 2*2*17/2*2*25( factorising it) =17/25 since both are equal neither of them are greater or smaller.

OpenStudy (anonymous):

ya they are equal